# Structured Unrestricted-Rank Matrices for Parameter Efficient Fine-tuning

### v2

@@ Abstract
Recent efforts to scale Transformer models have demonstrated rapid progress across a wide range of tasks [74]. However, fine-tuning these models for downstream tasks is expensive due to their large parameter counts. Parameter-efficient fine-tuning (PEFT) approaches have emerged as a viable alternative by allowing us to fine-tune models by updating only a small number of parameters. In this work, we propose a general framework for parameter efficient fine-tuning (PEFT), based on structured unrestricted-rank matrices (SURM) which can serve as a drop-in replacement for popular approaches such as Adapters and LoRA. Unlike other methods like LoRA, SURMs provides more flexibility in finding the right balance between compactness and expressiveness. This is achieved by using low displacement rank matrices (LDRMs), which hasn’t been used in this context before. SURMs remain competitive with baselines, often providing significant quality improvements while using a smaller parameter budget. SURMs achieve $\mathbf{5}$-$\mathbf{7}\%$ accuracy gains on various image classification tasks while replacing low-rank matrices in LoRA. It also results in up to $\mathbf{12x}$ reduction of the number of parameters in adapters (with virtually no loss in quality) on the GLUE benchmark.
**footnotetext: Equal Contribution
@@ 1 Introduction
We propose a novel paradigm of parameter efficient fine-tuning that leverages Structured Unrestricted-Rank matrices (or SURMs). SURMs provide efficiency gains obtained in prior works, but their less rigid structure is a gateway to quality improvements. We focus on the two sub-classes of SURMs: (1) Kronecker product of matrices [3] and (2) low displacement rank matrices (LDRMs) [64, 6, 56, 53, 70]. In this work, we extensively test the two following classes of LDRMs : circulant and Toeplitz matrices (see: Sec. 3). Our proposal is to perform PEFT by introducing additional weights parameterized as structured matrices. To summarize, our primary contributions are:
  1. 1.
    We demonstrate strong matrix approximation capabilities inherent in Low Displacement Rank Matrices, with a specific focus on circulant and Toeplitz matrices (Fig. 1 left, Fig. 4 and Sec. 4).
  2. 2.
    We propose the class of Structured Unrestricted-Rank matrices (SURMs) (Sec. 3), for parameter efficient fine-tuning of Transformers. SURMs include low-rank matrices used in LoRA, as special cases. To the best of our knowledge, we are the first to apply LDRMs in this context.
  3. 3.
    We achieve $\mathbf{5}$-$\mathbf{7}$accuracy gains over LoRA on a wide variety of image datasets as well as in low resource setting (VTAB-1k benchmark). In some cases SURMs outperform full fine-tuning, while using only $\mathbf{55}$k training parameters (Fig. 1, right).
  4. 4.
    We introduce a new class of adapter-layers using SURMs, achieving a $\mathbf{12x}$ reduction in parameters compared to adapters, with virtually no loss in quality on the GLUE benchmark (Sec. 6).
@@ 2 Related Work
With the introduction of BERT [20] and GPT-2 [8], Transformer models trained on general text corpora have revolutionized the field of machine learning (ML). Since then, these models have continued to increase in size, with open-source variants adopting various architectures. Examples include encoder-decoder models such as T5 [62] with up to 20B parameters [65], and a range of auto-regressive decoder models like Llama [67], Pythia [4], Mistral [28], among others, varying in size from a few million to 180B parameters [2]. These models have demonstrated general capabilities and can be easily adapted to downstream tasks by fine-tuning on task-specific data, resulting in state-of-the-art performance across a broad spectrum of downstream tasks. Due to the computational infeasibility of fine-tuning all the parameters of these models, in-context learning [8] and prompt engineering [11, 22] have emerged as attractive ways to adapt models to downstream tasks. However, such adaptation results depend heavily on the design of the input prompt and tend to vary greatly with small perturbations of the prompts [47].
@@
Consequently, many works have proposed various PEFT techniques. One of the earliest methods involves inserting the so-called adapter layers between existing layers in a neural network [25, 57]. A standard implementation of an adapter is an MLP with input-, output- and a smaller middle-layer, effectively encoded by two lower-rank matrices, and thus with a compact set of parameters. An extension of the adapter is Compacter [50], which uses Tucker decomposition to parameterize the adapter layers and weight-sharing to reduce the number of trainable parameters. Various modifications and extensions of the above methods have been proposed [52, 24, 31, 51, 63]. Another popular PEFT technique is differentiable prompt-tuning (DPT), which can be thought of as optimizing special tokens in the prompt [84]. However, these methods are limited by the sequence length of the underlying models. Even though DPT was originally developed for NLP, several works have extended it for computer vision tasks as well [79, 12, 27, 24].
@@
One of the most popular PEFT methods is Low Rank Adaptation [26], which imposes a low-rank constraint on the weight updates. The primary difference between the two approaches is that the learned LoRA weights can be merged with the frozen model weights during inference, thus not introducing any additional latency. This is not the case for the adapters. Given the popularity of LoRA, there has been a plethora of works on extending it to different contexts like long-range modeling [13], multi-tasking [10] or further improving its efficiency [19, 69, 45, 36, 30] among many others.
@@
Low-rank matrices are studied extensively within various ML problems [55, 83, 43, 59]. The research on low displacement rank matrices (LDRMs) for ML is more narrow  [85, 66, 40, 64, 15, 34, 60]. The application of Kronecker matrices in the context of PEFT is not new and has already been investigated in previous works [21, 24, 50], but in all the above cases the constituent matrices in the Kronecker product have low rank. In this work, we impose no such condition, while keeping a fixed parameter budget. To the best of our knowledge, we are the first to systematically explore the effectiveness of different structured matrices and introduce LDRMs for PEFT, extending beyond previous works.
@@
The rest of the paper is organized as follows: (a) We introduce the notion of Structured Unrestricted-Rank Matrices (SURM) that are used in this work (Sec 3), (b) We motivate the usage of SURM by empirically showing the approximation qualities of these matrices (Sec 4), (c) We use SURM as drop-in replacement for popular approaches such as Adapters and LoRA (Sec 5), (d) We validate our approach across a wide range of vision and NLP tasks (Sec 6).
3 Structured Unrestricted-Rank Matrices
@@
In this section, we will define the matrices that are used in this work. A Structured matrix is a generic term for a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ that can be represented by fewer than $mn$ parameters. This implies space (and often also time) complexity savings.
@@
A trivial example is a matrix of the form $\mathbf{W}=\mathbf{A}\mathbf{B}^{\top}\in\mathbb{R}^{m\times n}$ for $\mathbf{A}\in\mathbb{R}^{m\times r}$, $\mathbf{B}\in\mathbb{R}^{n\times r}$ with $r\ll\min(m,n)$. However, our main focus in this work is on those classes of structured matrices that are not restricted to be low-rank, which we refer to as Structured Unrestricted-Rank matrices (SURM).
@@
Low Displacement Rank Matrices :  Our first class of SURMs is low displacement rank matrices (LDRMs). A matrix $\mathbf{W}\in\mathbb{C}^{m\times n}$ is said to have $(\mathbf{A},\mathbf{B})$-displacement structure if:
@@
There exist several pairs of $(\mathbf{A},\mathbf{B})$ matrices $\mathbf{W}$ satisfying Eq. 1 with low-rank $\mathbf{F}$ that support fast (sub-quadratic) matrix-vector multiplication (as well as more efficient performance of other matrix-operations such as taking inversions, etc.). Some examples of such pairs include: $(\mathbf{Z},\mathbf{Z}),(\mathbf{Z},\mathbf{Z}^{\top}),(\mathbf{D}_{x},\mathbf{Z}^{\top}),(\mathbf{D}_{x},\mathbf{D}_{y})$ (for $x\neq y$). Here $\mathbf{Z}$ stands for the circulant-shift matrix and $\mathbf{D}_{z}$ is a diagonal matrix with nonzero entries equal to $z$. By choosing more complicated $(\mathbf{A},\mathbf{B})$-pairs, e.g. using general Jordan form matrices, one can consider more unstructured $\mathbf{W}$, yet still possessing compact representations and supporting efficient matrix operations [54, 64]. In this paper, we focus on classic LDRMs: circulant and toeplitz matrices described below.
@@
    Circulant Matrices : A circulant matrix $\mathbf{C}\in\mathbb{C}^{m\times n}$ can be parameterized by its first row, with next row obtained from the previous one by the right circulant shift. A schematic visualization is provided in Fig 2 (a). Circulant matrices can be trivially encoded in $O(n)$ space and support fast $O((n+m)\log(n+m))$ matrix-vector multiplication via Fast Fourier Transform (FFT).
@@
    Toeplitz Matrices : A Toeplitz matrix $\mathbf{T}\in\mathbb{C}^{m\times n}$, shown in Fig 2 (b), are constant along each diagonal. They can be parameterized using only their first row and column and can be encoded in $O(n+m)$ space. Like circulant matrices, they support fast $O((n+m)\log(n+m))$ matrix-vector multiplication via FFT.
Figure 2: A Schematic Diagram of Circulant, Toeplitz and Kronecker product of 2 matrices.
@@
Kronecker Product of Matrices. These matrices are not necessarily LDRMs, but they also enjoy low storage complexity and admit efficient matrix-vector multiplication. A schematic diagram showing the Kronecker product of two matrices is provided in Fig 2 (c). In our experiments, we apply Kronecker product of two matrices. We provide more details about these matrices in Appendix A.2.
4 LDR-SURMs as General Approximators
In this section, we motivate the usage of structured unrestricted-rank matrices (SURMs) for parameter efficient fine-tuning by showcasing their role as approximators of various classes of matrices. Without loss of generality, we assume that all our matrices have real entries and $n=m$ ( square matrices). We denote the rank of $\mathbf{F}$ from Eq. 1 as $r$.
It is shown in [64] that a large class of LDRMs, including (a) circulant, Toeplitz and inverse-Toeplitz matrices as well as: (b) linear combinations of the products of the form: $\mathbf{M}_{1}\cdot...\cdot\mathbf{M}_{t}$ for $r\geq 2t$ and where each $\mathbf{M}_{i}$ is a Toeplitz matrix or its inverse, can be parameterized as follows for $\mathbf{g}_{1},...,\mathbf{g}_{r},\mathbf{h}_{1},...,\mathbf{h}_{r}\in\mathbb{R}^{n}$:
$\mathbf{W}(\mathbf{G},\mathbf{H})=\sum_{i=1}^{r}\mathbf{Z}_{1}(\mathbf{g}_{i})\mathbf{Z}_{-1}(\mathbf{h}_{i}).$ (2)
Here: $\mathbf{Z}_{f}(\mathbf{v})$, for $f\in\mathbb{R}$ and $v\in\mathbb{R}^{n}$, is defined as:
$\scriptstyle{{\displaystyle\mathbf{Z}_{f}(\mathbf{v})={\begin{bmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}v_{0}}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}fv_{n-1}}&\cdots&{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}fv_{1}}\\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}v_{1}}&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}v_{0}}&\cdots&{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}fv_{2}}\\ \vdots&\vdots&\vdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}fv_{n-1}}\\ {\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}v_{n-1}}&\cdots&{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}v_{1}}&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}v_{0}}\\ \end{bmatrix}}}}$ (3)
Besides, $\mathbf{F}$ can be decomposed as follows: $\mathbf{F}=\mathbf{G}\mathbf{H}^{\top}$ for $\mathbf{G}=[\mathbf{g}_{1},...,\mathbf{g}_{r}],\mathbf{H}=[\mathbf{h}_{1},...,\mathbf{h}_{r}]\in\mathbb{R}^{n\times r}$. One can think about rank $r$ of $\mathbf{F}$ of controlling how “structured” $\mathbf{W}$ is.
We conduct the approximation in two settings: (a) comparing the approximation qualities of $\mathbf{W}(\mathbf{G},\mathbf{H})$ with circulant and Toeplitz matrices, and (b) comparing circulant and Toeplitz matrices with low-rank matrices.
4.1 $\mathbf{W}(\mathbf{G},\mathbf{H})$ versus circulant and Toeplitz matrices
We test the approximation capabilities of matrices $\mathbf{W}(\mathbf{G},\mathbf{H})$ from Eq. 2 as well as circulant and Toeplitz matrices to approximate three classes of matrices: (a) random, (b) near-low-rank, and (c) near-low-intrinsic-rank. Matrices $\mathbf{M}$ from all the classes are taken from $\mathbb{R}^{100\times 100}$. Detailed description of these matrices are presented in Appendix A.5.
The approximating structured matrix is obtained via gradient descent on the loss function $\|\mathbf{A}-\mathbf{M}\|_{\mathrm{F}}^{2}$, where $\mathbf{A}$ and $\mathbf{M}$ are: the approximating and ground truth matrix and $\|\cdot\|_{\mathrm{F}}$ is the Frobenius norm. However, for circulant and Toeplitz matrices closed-form formula for the optimal approximating matrix is available (see Fig 3 and Appendix A.4).
$\mathbf{W}(\mathbf{G},\mathbf{H})$ with different $r$ is used to approximate random matrices (Fig 5 (Top Left)). While the best approximations are achieved for larger values of $r$ (specifically, $r=20$), it’s interesting to note that the final accuracy does not exhibit a monotonic increase with $r$. Consequently, for the second class of matrices, which are easier to approximate and are close to low-rank, we experiment with smaller values of $r$ (Fig 5 (Left Column Middle and Bottom)). In this case, the three top-performing approximators were trained with $r=2$, $r=1$, and $r=4$. These results clearly show that for more structured ground truth matrices (even though not necessarily low-rank !), LDRMs with very low rank of the corresponding $\mathbf{F}$ are sufficient.
For that reason, we focus our attention on circulant and Toeplitz matrices. In the next three plots in Fig. 5 (Right Column), we compare their approximation capabilities for near-to-low-rank matrices as well as low-intrinsic-rank matrices. We conclude that the gains coming from Toeplitz matrices (that use two times as many parameters as circulant matrices) are negligible and present only in the near-low-rank case. For the low-intrinsic case, circulant matrices outperform Toeplitz ones.
4.2 Low-rank versus Circulant and Toeplitz matrices

### v3

@@ Abstract
Recent efforts to scale Transformer models have been successful across a wide range of tasks [77]. However, fine-tuning these models for downstream tasks can be expensive, as it requires updating a large number of parameters in the Transformer model. Parameter-efficient fine-tuning (PEFT) approaches have emerged as a viable alternative that allow us to fine-tune models by updating only a small number of parameters. In this work, we propose a general framework for parameter efficient fine-tuning using structured unrestricted-rank matrices (SURM), which can serve as a drop-in replacement for popular approaches such as Adapters and LoRA. Unlike other methods like LoRA, SURMs provides more flexibility in finding the right balance between compactness and expressiveness. This is achieved by using low displacement rank matrices (LDRMs), which has not been used in this context before. SURMs remain competitive with baselines, often providing significant quality improvements while using a smaller parameter budget. SURMs achieve $\mathbf{5}$-$\mathbf{7}$accuracy gains on various image classification tasks while replacing low-rank matrices in LoRA. It also results in up to $\mathbf{12x}$ reduction of the number of parameters in adapters (with virtually no loss in quality) on the GLUE benchmark.
@@ 1 Introduction
We propose a novel paradigm of parameter efficient fine-tuning that leverages Structured Unrestricted-Rank matrices (or SURMs). SURMs provide similar efficiency gains as previous works in efficient fine-tuning, but their more expressive structure paves the way for quality improvements. In this work, we propose to perform parameter-efficient fine-tuning by parameterizing learnable weights as structured matrices. We focus on the two sub-classes of SURMs: (1) Kronecker product of matrices [3] and (2) low displacement rank matrices (LDRMs) [66, 6, 58, 54, 73]. To summarize, our primary contributions are:
  • We propose the class of Structured Unrestricted-Rank matrices (SURMs) (Section 3), for parameter efficient fine-tuning of Transformers. SURMs include low-rank matrices used in LoRA, as special cases. To the best of our knowledge, we are the first to apply LDRMs in this context.
  • We demonstrate strong matrix approximation capabilities inherent in Low Displacement Rank Matrices, with a specific focus on circulant and Toeplitz matrices (Section 4).
  • We introduce a new class of adapter-layers using SURMs, achieving a $\mathbf{12x}$ reduction in parameters compared to adapters, with virtually no loss in quality on the GLUE benchmark (Section 6).
  • We achieve $\mathbf{5}$-$\mathbf{7}$accuracy gains over LoRA on a wide variety of image datasets as well as in low resource setting (VTAB-1k benchmark). In some cases SURMs outperform full fine-tuning, while using only $\mathbf{55}$k training parameters (as shown in Fig. 1 (right)).
@@ 2 Related Work
With the introduction of BERT [20] and GPT-2 [8], Transformer models trained on general text corpora have revolutionized the field of machine learning (ML). Since then, these models have continued to increase in size, with open-source variants adopting various architectures. Examples include encoder-decoder models such as T5 [64] with up to 20B parameters [67], and a range of auto-regressive decoder models like Llama [69], Pythia [4], Mistral [28], among others, varying in size from a few million to 180B parameters [2]. These models can be easily adapted to downstream tasks by fine-tuning on task-specific data, resulting in state-of-the-art performance across a broad spectrum of downstream tasks. Due to the computational infeasibility of fine-tuning all the parameters of these models, in-context learning [8] and prompt engineering [11, 22] have emerged as attractive alternatives to adapt models to downstream tasks. However, such adaptation results depend heavily on the design of the input prompt and tend to vary greatly with small perturbations of the prompts [48].
@@
Consequently, many works have proposed various PEFT techniques. One of the earliest methods involves inserting the so-called adapter layers between existing layers in a neural network [25, 59]. An adapter is typically an MLP with input, output, and a smaller middle layer, encoded by two low-rank matrices, making it compact in terms of parameters. An extension of the adapter is Compacter [51], which uses Tucker decomposition to parameterize the adapter layers and weight-sharing to reduce the number of trainable parameters. Various modifications and extensions of the above methods have been proposed [53, 24, 32, 52, 65]. Another popular PEFT technique is differentiable prompt-tuning (DPT), which can be thought of as optimizing special tokens in the prompt [88]. However, these methods are limited by the sequence length of the underlying models. Even though DPT was originally developed for NLP, several works have extended it for computer vision tasks as well [83, 12, 27, 24].
@@
One of the most popular PEFT methods is Low-Rank Adaptation (LoRA) [26], which imposes a low-rank constraint on the weight updates. The main difference between adapters and LoRA is that the learned LoRA weights can be merged with the frozen model weights during inference without adding any latency. Given the popularity of LoRA, there have been many works on extending it to different contexts like long-range modeling [13], multi-tasking [10] or improving its efficiency [19, 71, 46, 37, 31] among many others.
@@
In general, low-rank matrices are studied extensively in various ML applications [57, 87, 44, 61]. The research on low displacement rank matrices (LDRMs) for ML is more narrow  [89, 68, 41, 66, 15, 35, 62]. Although Kronecker matrices (a class of LDRMs) have been explored in the context of LDRMs [21, 24, 51], the constituent matrices in the Kronecker product have low rank even in these work. In this work, we use a fixed parameter budget but do not impose any rank-based condition. To the best of our knowledge, we are the first to systematically explore the effectiveness of different structured matrices and introduce LDRMs for parameter-efficient fine-tuning.
@@
The rest of the paper is organized as follows: (a) We introduce the notion of Structured Unrestricted-Rank Matrices (SURM) that are used in this work (Section 3), (b) We motivate the usage of SURM by empirically showing the approximation qualities of these matrices (Section 4), (c) We use SURM as drop-in replacement for popular approaches such as Adapters and LoRA (Section 5), (d) We validate our approach across a wide range of vision and NLP tasks (Section 6).
3 Structured Unrestricted-Rank Matrices (SURM)
@@
In this section, we will define the matrices that are used for parameter efficient fine-tuning. First, we define the concept of a structured matrix, which is a generic term for a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ that can be represented by fewer than $mn$ parameters. These matrices are useful because they reduce both space and time complexity when performing matrix multiplications.
@@
A simple example of a structured matrix is a low rank matrix of the form $\mathbf{W}=\mathbf{A}\mathbf{B}^{\top}\in\mathbb{R}^{m\times n}$, where $\mathbf{A}\in\mathbb{R}^{m\times r}$, $\mathbf{B}\in\mathbb{R}^{n\times r}$ with $r\ll\min(m,n)$. In this work, our main focus is on those classes of structured matrices that are not restricted to be low-rank, which we refer to as Structured Unrestricted-Rank Matrices (SURM). Next, we present two classes of SURM matrices that we use for parameter efficient fine-tuning.
@@
Low Displacement Rank Matrices. Our first class of SURMs is low displacement rank matrices (LDRMs). A matrix $\mathbf{W}\in\mathbb{C}^{m\times n}$ is said to have $(\mathbf{A},\mathbf{B})$-displacement structure if:
@@
For a given $\mathbf{W}$, there can exist several pairs of $(\mathbf{A},\mathbf{B})$ matrices satisfying Eq. 1 that produce a low-rank matrix, $\mathbf{F}$. Some examples of such $(\mathbf{A},\mathbf{B})$ pairs include: $(\mathbf{Z},\mathbf{Z}),(\mathbf{Z},\mathbf{Z}^{\top}),(\mathbf{D}_{x},\mathbf{Z}^{\top}),(\mathbf{D}_{x},\mathbf{D}_{y})$ (for $x\neq y$). Here $\mathbf{Z}$ is a circulant-shift matrix and $\mathbf{D}_{z}$ is a diagonal matrix with nonzero entries equal to $z$. Low displacement rank matrices ($\mathbf{W}$ in Eq. 1) enable fast (sub-quadratic) matrix-vector multiplication and enhance the efficiency of other matrix operations, such as inversion. By selecting more complex $(\mathbf{A},\mathbf{B})$ pairs, such as those involving general Jordan form matrices, it is possible to consider more unstructured $\mathbf{W}$ that still have compact representations and support efficient matrix operations [56, 66]. In this paper, we focus on classic low displacement rank matrices: circulant and toeplitz matrices, which are described below.
@@
    Circulant Matrices: A circulant matrix $\mathbf{C}\in\mathbb{C}^{m\times n}$ can be parameterized by its first row. The following rows are obtained from the previous one by applying a right circulant shift. A schematic visualization of a circulant matrix is shown in Fig 2 (a). Since we only need to store the first row, circulant matrices can be trivially encoded in $O(n)$ space. They also support fast matrix-vector multiplication in $O((n+m)\log(n+m))$ time using Fast Fourier Transform (FFT) [55].
@@
    Toeplitz Matrices: A toeplitz matrix $\mathbf{T}\in\mathbb{C}^{m\times n}$ is a matrix whose entries are constant along each diagonal. A schematic visualization of a toeplitz matrix is shown in Fig 2 (b)). They can be parameterized using only their first row and column, allowing them to be encoded in $O(n+m)$ space. Similar to circulant matrices, they support fast $O((n+m)\log(n+m))$ matrix-vector multiplication via FFT.
Figure 2: A schematic diagram to illustrate the structure (a) Circulant, (b) Toeplitz, and (c) Kronecker product of two matrices $\mathbf{A}$ and $\mathbf{B}$.
Kronecker Product of Matrices. Kronecker products are another class of structured unrestricted rank matrices that have low storage complexity and admit efficient matrix-vector multiplication. These matrices are obtained using a Kronecker product $\mathbf{A}\otimes\mathbf{B}$ of two matrices $\mathbf{A}$ and $\mathbf{B}$, as shown in Fig 2 (c). We provide more details about these matrices in Appendix A.3.
4 LDR-SURMs as General Approximators
In this section, we motivate the usage of structured unrestricted-rank matrices (SURMs) for parameter-efficient fine-tuning. In general, the parameter updates $\Delta\mathbf{W}$ can be arbitrary matrices, and an effective parameterization of $\Delta\mathbf{W}$ should be sufficiently expressive to approximate them. Since we use structured update rank matrices (SURMs) to parameterize $\Delta\mathbf{W}$, we demonstrate that SURMs can approximate various classes of matrices. Without loss of generality, in this section, we assume that all our matrices have real entries and that weight matrices are square ($n=m$).
[…]
$\mathbf{W}(\mathbf{G},\mathbf{H})=\sum_{i=1}^{r}\mathbf{Z}_{1}(\mathbf{g}_{i})\mathbf{Z}_{-1}(\mathbf{h}_{i}),$ (2)
where $\mathbf{G}=[\mathbf{g}_{1},...,\mathbf{g}_{r}],\mathbf{H}=[\mathbf{h}_{1},...,\mathbf{h}_{r}]\in\mathbb{R}^{n\times r}$, and $\mathbf{Z}_{f}(\mathbf{v})$ (for any $f\in\mathbb{R}$, $v\in\mathbb{R}^{n}$) is defined as:
${}\scriptstyle{{\displaystyle\mathbf{Z}_{f}(\mathbf{v})={\begin{bmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}v_{0}}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}fv_{n-1}}&\cdots&{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}fv_{1}}\\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}v_{1}}&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}v_{0}}&\cdots&{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}fv_{2}}\\ \vdots&\vdots&\vdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}fv_{n-1}}\\ {\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}v_{n-1}}&\cdots&{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}v_{1}}&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}v_{0}}\\ \end{bmatrix}}}}.$ (3)
When $f=1$, $\mathbf{Z}_{f}(v)$ is a circulant matrix and when $f=-1$, we refer to $\mathbf{Z}_{f}(v)$ as a skew-circulant matrix. Moreover, $\mathbf{F}$ can be decomposed as follows: $\mathbf{F}=\mathbf{G}\mathbf{H}^{\top}$ for $\mathbf{G}=[\mathbf{g}_{1},...,\mathbf{g}_{r}],\mathbf{H}=[\mathbf{h}_{1},...,\mathbf{h}_{r}]\in\mathbb{R}^{n\times r}$. One can think about rank $r$ of $\mathbf{F}$ of controlling how “structured” $\mathbf{W}$ is.
@@ 4 LDR-SURMs as General Approximators
From the above result, we see that $\mathbf{W}(\bf G,\bf H)$ is the most expressive parameterization among the ones discussed so far. To understand how they fare with SURMs in practice, we evaluate their approximation qualities in two settings: (a) comparing $\mathbf{W}(\mathbf{G},\mathbf{H})$ with circulant and Toeplitz matrices, and (b) comparing circulant and Toeplitz matrices with low-rank matrices.
4.1 Comparing $\mathbf{W}(\mathbf{G},\mathbf{H})$ with Circulant and Toeplitz Matrices
We test the approximation capabilities of matrices $\mathbf{W}(\mathbf{G},\mathbf{H})$ (Eq. 2) and compare it with popular classes of SURMs: circulant and Toeplitz matrices (Section 3). Specifically, we use these structured matrices to approximate three broad classes of matrices: (a) random, (b) near-low-rank, and (c) near-low-intrinsic-rank. We denote the ground-truth matrix that we try to approximate as $\mathbf{M}\in\mathbb{R}^{100\times 100}$ and parameterized structured matrix as $\mathbf{A}$ (see more details about the setup in Appendix A.7). For all matrices, we obtain the parameters of $\mathbf{A}$ using gradient descent on the loss function: $\|\mathbf{A}-\mathbf{M}\|_{\mathrm{F}}^{2}$.  [^2] Please note that for circulant and Toeplitz matrices, it is also possible to obtain a closed-form solution for the matrix (see Fig 3 & Appendix A.6).